OpenStudy (anonymous):

A printer will produce 80 wedding invitations for $210. The price to produce 120 invitations is $290. The printer uses a linear function to determine the price for producing different amounts of invitations. How much will the printer charge to produce 60 invitations? $120.00 $145.00 $157.50 $170.00

OpenStudy (mathmate):

Hint: Use the equation y=mx+b where m=(y2-y1)/(x2-x1) 

OpenStudy (anonymous):

i did m=(y2-y1)/(x2-x1) and got 2

OpenStudy (anonymous):

so y = 2x + b

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

Good! Now can you find b, knowing that when x=80, y=210?